OpenStudy (anonymous):

3a + 2[2 + 3(a − 1)] = 2(3a + 4) im not getting this. i keep trying the distributive properties, but im not getting the correct answer/ they give me the answers and i have to show the work...answer is 10/3

OpenStudy (anonymous):

3a+2(2+3a-3)=2(3a+4) 3a+4+6a-6=6a+8 9a-2=6a+8 9a-6a=8+2 3a=10 a=10/3